OpenStudy (anonymous):

Juan's gas tank is 1/3 full. After he buys 13 gallons of gas, it is 5/6 full. How many gallons can Juan's tank hold?

OpenStudy (anonymous):

@ksanka Is this one 26

OpenStudy (anonymous):

1/3 -5/6=-1/2

OpenStudy (anonymous):

2*13=26

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

I got that one correct

OpenStudy (anonymous):

5/6 - 1/3 = 1/2 1/2 of tank is 13 gallons. So 13*2=26. Everything is ok except for first operation.

OpenStudy (anonymous):

I have another one

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

The container that holds the water for the football team is 1/4 full. After pouring in 10 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

2/3 - 1/4 = (8-3)/12=5/12 5/12 of container are 10 gallons. 10/(5/12))=10*12/5=24